OpenStudy (anonymous):

WILL GIVE METAL!! AND FAN!! A television network is about to telecast a new television show. Before a show is launched, the network airs a pilot episode and receives a report assessing favorable or unfavorable viewer response. In the past, 60% of the network's shows have received a favorable response from viewers, and 40% have received an unfavorable response. If 50% of the network’s shows have received a favorable response and have been successful, and 30% of the network’s shows have received

OpenStudy (anonymous):

an unfavorable response and have been successful, what is the probability that this new show will be successful if it receives a favorable response?

OpenStudy (anonymous):

put A is the event the show is successful B is the event the show receives a favorable response and write the probability of A given B as \[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

both the numbers on the right are given to you \(P(B)\) is the probability that a show receives a favorable response. you are told it is \(60\%=.6\)

OpenStudy (anonymous):

\[P(A\cap B)\] is the probability that it is successful AND receives a favorable response you are told this is \(50\%=.5\) in this line 50% of the network’s shows have received a favorable response and have been successful

OpenStudy (anonymous):

your last job is to compute \[\frac{P(A\cap B)}{P(B)}\] which in this case is \(\frac{.5}{.6}=\frac{5}{6}\)

OpenStudy (anonymous):

the options are 0.41 0.53 0.67 0.70 0.83

OpenStudy (anonymous):

83 :)